how do you create a table of six values that does not represent a linear relation. All the y-values must increase in values as x increase in value?
Answer by Alan3354(69443)

how do you create a table of six values that does not represent a linear relation. All the y-values must increase in values as x increase in value?
-----------
Linear is 1st order, an exponent of 1.
Make it any other exponent.
eg, y = x^1.2 or y = x^0.5